# Entanglement, Spin Squeezing and Quantum Sensing in a Spin-5/2 Heisenberg Molecular Iron(III) Triangle

### v2

@@ III Quantum entanglement
The bipartite negativity Vidal2002 represents one of the most commonly used measures of the pairwise entanglement in the quantum information theory, which quantifies the amount of nonclassical correlations between two subsystems of a composite quantum system. Owing to the $\mathrm{C}_{3}$ symmetry of the trinuclear complex Fe3, all three reduced density matrices of the spin-5/2 Heisenberg triangle obtained after tracing out degrees of freedom of the third spin ${\rho}_{ij}=\mbox{Tr}_{k}\rho$ are completely identical ${\rho}_{12}={\rho}_{23}={\rho}_{13}\equiv{\rho}_{\mathrm{Bip}}$. Consequently, all three associated bipartite negativities ${N}_{12}={N}_{23}={N}_{13}\equiv{N}_{\mathrm{Bip}}$ will become identical as well. To quantify a strength of the bipartite entanglement between any pair of spin-5/2 magnetic ions iron(III) of the trinuclear complex Fe3 one can therefore utilize the formula put forward by Vidal and Werner Vidal2002 , which is defined according to the Peres–Horodecki separability criterion Peres1996 ; Horodecki1996 as a sum of absolute values of negative eigenvalues of a partially transposed reduced density matrix $\rho_{\mathrm{Bip}}^{\mathrm{T}}$
@@
Here, $\lambda_{i}$ denote the eigenvalues of the reduced density matrix $\rho_{\mathrm{Bip}}^{\mathrm{T}}$ that is partially transposed with respect to one of two remaining spins. The reduced density matrix $\rho_{\mathrm{Bip}}$ and its transpose $\rho_{\mathrm{Bip}}^{\mathrm{T}}$ along with the eigenvalues $\lambda_{i}$ are obtained using exact numerical methods implemented in the QuTip package Johansson2012 .
Figure 3: (a) Magnetic-field dependencies of the bipartite negativity ${N}_{\mathrm{Bip}}$ for a few different values of temperature; (b) temperature dependencies of the bipartite negativity ${N}_{\mathrm{Bip}}$ for a few selected values of the magnetic field. In both panels we present the theoretical prediction of the bipartite negativity ${N}_{\mathrm{Bip}}$ for any pair of high-spin iron(III) magnetic ions within the molecular complex Fe3 based on the spin-5/2 Heisenberg triangle, which is given by the Hamiltonian (1) with the coupling constant $J=12.56\,\mathrm{cm}^{-1}$ and the gyromagnetic factor $g=2.0$ as determined from the previous experiments Plass2023 .
@@
To determine whether all three spin-5/2 iron(III) magnetic ions in the trinuclear complex Fe3 are involved in multipartite quantum correlations, we further examine the tripartite entanglement within the respective spin-5/2 Heisenberg triangle. The tripartite quantum entanglement shared among all three spins is captured by the quantity genuine tripartite negativity ${N}_{\mathrm{Trip}}$, which is defined as the geometric mean of the bipartite negativities calculated for all three possible bipartitions of the spin-5/2 Heisenberg triangle Sabin2008
@@
Figure 4: (a) Magnetic-field dependencies of the tripartite negativity ${N}_{\mathrm{Trip}}$ for a few different values of temperature; (b) temperature dependencies of the tripartite negativity ${N}_{\mathrm{Trip}}$ for a few selected values of the magnetic field. In both panels we present the theoretical prediction of the tripartite negativity ${N}_{\mathrm{Trip}}$ for any pair of high-spin iron(III) magnetic ions within the molecular complex Fe3 based on the spin-5/2 Heisenberg triangle, which is given by the Hamiltonian (1) with the coupling constant $J=12.56\,\mathrm{cm}^{-1}$ and the gyromagnetic factor $g=2.0$ as determined from the previous experiments Plass2023 .
@@
We now proceed with a theoretical examination of the genuine tripartite negativity ${N}_{\mathrm{Trip}}$ of the trinuclear complex Fe3 as shown in Fig. 4. We note that the QuTip package Johansson2012 is used here to numerically obtain the partially transposed density matrix $\rho_{\mathrm{ABC}}^{\mathrm{T}_{\mathrm{A}}}$ and its eigenvalues. It is evident from Fig. 4(a) that the behavior of the genuine tripartite negativity closely resembles at low temperatures that of its bipartite counterpart depicted in Fig. 3(a). Specifically, the genuine tripartite negativity ${N}_{\mathrm{Trip}}$ repeatedly exhibits unconventional step-like changes including a series of plateaus and sudden downturns as the magnetic field increases. As the temperature rises, this stepwise behavior of the genuine tripartite negativity ${N}_{\mathrm{Trip}}$ is gradually fading and transitioning into a smooth monotonic decrease with increasing magnetic field (see the lines for $T\gtrsim 10$ K in Fig. 4(a)). Additional insights can be drawn from the temperature dependencies of the genuine tripartite negativity in Fig. 4(b), which reveals that the molecular complex Fe3 displays a high degree of tripartite entanglement at low enough temperatures $T\lesssim 10$ K and magnetic fields $B\approx 10\,\mathrm{T}$. Under these conditions, the eigenstate $|1/2\rangle$ with the lowest possible spin value dominates the overall mixed state. Notably, the tripartite negativity gradually decreases as the temperature increases and eventually vanishes at a significantly higher threshold temperature $T\approx 70\,\mathrm{K}$ compared to the threshold temperature of the bipartite entanglement when exceeding it more than twice. This finding suggests that the trinuclear complex Fe3 remains subject to multipartite entanglement even at elevated temperatures and magnetic fields.
IV Spin Squeezing
The concept of spin squeezing, which was introduced in the seminal works of Kitagawa and Ueda kitagawa1991 ; kitagawa1993 , is a powerful technique at the forefront of precision measurement with potential applications in various fields such as quantum metrology, computation and simulation ma2011 ; gross2012 . In spin squeezing, the focus is on ensembles of correlated spins, i.e., collective spins wineland1992 ; wineland1994 , which are often referred to as pseudospins. These ensembles can achieve relatively large collective spin values hazzard2013 ; bilitewski2021 ; Tscherbul2023 .
The spin squeezing phenomenon is characterized by the spin squeezing parameter $\xi^{2}$, which quantifies the reduction of spin fluctuations in one spatial direction at the expense of increased fluctuations in another direction(s). Mathematically, the spin squeezing parameter $\xi^{2}$ is defined as follows:
$\xi^{2}=\frac{2(\Delta J_{\mathrm{min}})^{2}}{|\langle\hat{\bm{J}}_{n}\rangle|},$ (6)
where $\Delta J_{\mathrm{min}}$ represents the minimum uncertainty (or fluctuation) in the measurement of the collective spin component along a specific direction and $\langle\hat{\bm{J}}_{n}\rangle$ is the expectation value of the collective spin operator along another direction orthogonal to $\Delta J_{\mathrm{min}}$. The spatial components of the total spin angular momentum $\hat{\bm{J}}$ of the spin-$5/2$ Heisenberg triangle are given by:
$\hat{J}_{\alpha}=\sum\limits_{i=1}^{3}\hat{S}_{i}^{\alpha}\quad\mathrm{with}\quad\alpha=\{x,y,z,\pm\}.$ (7)
By adapting the formula derived in Refs. Cheraghi2022 ; Mahdavifar2023 ; Mahdavifar2024 one gets the following definition of the spin squeezing parameter of the spin-5/2 Heisenberg triangle composed from three identical spin-5/2 entities:
$\xi^{2}=\frac{2}{3}\left[\langle J_{x}^{2}+J_{y}^{2}\rangle-\sqrt{\langle J_{x}^{2}-J_{y}^{2}\rangle^{2}+\langle J_{x}J_{y}+J_{y}J_{x}\rangle^{2}}\right],$ (8)
A value of the spin squeezing parameter $\xi^{2}<1$ indicates that the spin state is squeezed. This means that the quantum spin fluctuations in one direction are reduced below the limit set by the uncertainty principle at the cost of increased fluctuations in another direction(s). This reduction in uncertainty can enhance sensitivity in measurements of the collective spin what makes the spin squeezing a valuable tool in precision measurements and quantum technologies.
Figure 5: (a) Spin squeezing parameter $\xi^{2}$ as a function of the magnetic field at a few different temperatures; (b) temperature variations of the spin squeezing parameter $\xi^{2}$ for several values of the magnetic field. In both panels we present theoretical predictions for the spin squeezing parameter $\xi^{2}$ of the molecular complex Fe3 as modeled by the spin-5/2 Heisenberg triangle (1) with the coupling constant $J=12.56\,\mathrm{cm}^{-1}$ and the gyromagnetic factor $g=2.0$ Plass2023 . A dotted line indicates $\xi^{2}=1$ showcases a coherent state. Below this line, the state is squeezed, whereas above it the state is unsqueezed.
Let us at first explore the dependence of the spin squeezing parameter on the magnetic field within the thermal state of the spin-5/2 Heisenberg equilateral triangle at various temperatures. It follows from Fig. 5(a) that the molecular compound Fe3 can exist, at sufficiently low temperatures $T\lesssim 15$ K, either in a squeezed state ($\xi_{s}^{2}<1$), a coherent state ($\xi_{s}^{2}=1$), or an unsqueezed state ($\xi_{s}^{2}>1$). The ground state $|\frac{3}{2}\rangle$, which emerges in a range of moderate magnetic fields $21\,\mathrm{T}\lesssim B\lesssim 34\,\mathrm{T}$ and gives rise to the intermediate 1/5-plateau in a zero-temperature magnetization curve (see a dotted line in Fig. 2), can be identified as a coherent state with $\xi_{s}^{2}=1$. For completeness, Fig. 5(b) displays the squeezing parameter $\xi_{s}^{2}$ as a function of temperature $T$ for various values of the magnetic field. It is evident that the molecular complex Fe3 may be at sufficiently low magnetic fields $B\lesssim 25\,\mathrm{T}$ in a spin-squeezed state with $\xi_{s}^{2}<1$. However, the trinuclear molecular compound Fe3 surpasses the coherent limit $\xi_{s}^{2}=1$ and enters a fully unsqueezed regime as the temperature increases. The most intriguing behavior of the spin squeezing parameter can be thus observed for the magnetic field $B\gtrsim 20\,\mathrm{T}$, where the system starts in a coherent state at low temperatures, then it proceeds to a squeezed state when reaching a minimum value of $\xi_{s}^{2}$ indicating maximum spin squeezing before it ultimately becomes unsqueezed when crossing the coherent limit as the temperature increases.
This intricate behavior highlights the delicate interplay between the magnetic field and temperature in determining the spin squeezing characteristics of the molecular complex Fe3. Understanding this interplay is crucial for optimizing conditions in applications that require precise control of quantum states. The transition from a coherent state to a spin squeezed state accompanied by the enhancement of spin squeezing is essential for applications in quantum information processing and quantum metrology. These findings highlight the potential for optimizing quantum states in the molecular compound Fe3 by carefully tuning both the magnetic field and temperature.
V Quantum Sensing
@@
To adapt the quantum sensing protocol suggested in Ref. Montenegro2022 for a real magnetic compound, let us consider the molecular complex Fe3 consisting of three interacting spin-5/2 iron(III) magnetic ions as a probe designed for sensing a local magnetic field $B_{x}$ acting on one iron(III) magnetic ion via measuring either of the remaining two iron(III) magnetic ions. The Hamiltonian of the spin-5/2 Heisenberg triangle with a local magnetic field acting only on a single spin, which may serve as the probe, can be updated to the following form:
@@
It has been demonstrated in Ref. Nepomechie that the Dicke states $|D_{k}\rangle$ of three qudits can be straightforwardly obtained by applying $k$ times the total spin-raising operator $J_{+}$ on the basis state with the lowest total spin angular momentum $|-\frac{5}{2},-\frac{5}{2},-\frac{5}{2}\rangle$. More specifically, the Dicke states $|D_{k}\rangle$ of three spin-5/2 qudits can be defined as follows:
@@
To demonstrate how the molecular compound Fe3 can facilitate the implementation of a quantum sensing protocol described in Ref. Montenegro2022 , we specifically consider a local magnetic field acting on the first spin of the spin-5/2 Heisenberg triangle with the readout performed sequentially on either of the remaining two spins.
Figure 6: The time evolution of the local magnetization of the first and third spin of the spin-5/2 Heisenberg triangle with the coupling constant $J=12.56\,\mathrm{cm}^{-1}$ and the gyromagnetic factor $g=2.0$ adjusted for a theoretical modeling of the molecular complex Fe3 Plass2023 by considering four different initial Dicke states: (a)–(b) $|D_{0}\rangle$; (c)–(d) $|D_{1}\rangle$; (e)–(f) $|D_{2}\rangle$; (g)–(h) $|D_{3}\rangle$. In the left panel we suppose a lower magnetic field $B_{x}=1\,\mathrm{T}$, while in the right panel we consider a higher magnetic field $B_{x}=5\,\mathrm{T}$. Both spins exhibit almost identical spin dynamics.
@@
After initializing, the state of the spin-5/2 Heisenberg triangle evolves under the influence of the Hamiltonian $\hat{H}^{\prime}$ and the dynamics of local magnetization $\langle S^{z}_{1}\rangle$ and $\langle S^{z}_{3}\rangle$ depends on the magnetic field. Fig. 6 illustrates the time evolution of the magnetization dynamics for different initial states under two different magnetic fields $B_{x}=1\,\mathrm{T}$ and $B_{x}=5\,\mathrm{T}$. Each panel compares the dynamical behavior of the readout spin $\langle S_{3}^{z}\rangle$ (pale lines) and the sensor spin $\langle S_{1}^{z}\rangle$ (blue dotted lines) for two different magnetic fields. Figs. 6(a) and 6(b) illustrate the time evolution of magnetization for the initial state $|D_{0}\rangle$ with character of the fully polarized ferromagnetic state. In Fig. 6(a), the sensor spin shows at a relatively small magnetic field $B_{x}=1\,\mathrm{T}$ oscillatory harmonic dependence and the readout spin closely follows this dependence. Fig. 6(b) shows the time evolution of the same quantities at the higher magnetic field $B_{x}=5\,\mathrm{T}$, for which the readout spin still mimics time evolution of the sensor spin but the oscillations are of higher frequency and less regular with gradually decreasing (and increasing) amplitude over the time. This scenario suggests more complex dynamical behavior as the magnetic field strengthens though the readout spin still continues to follow the sensor spin.
@@
Figs. 6(c)-6(h) depict the time evolution of magnetization for more complex initial Dicke states being subject to quantum superposition of states. At sufficiently small magnetic fields, the sensor and readout spins exhibit regular synchronized harmonic oscillations in time as demonstrated in Figs. 6(c), 6(e) and 6(g) for the particular case with $B_{x}=1\,\mathrm{T}$ and a few different Dicke states $|D_{1}\rangle$, $|D_{2}\rangle$, and $|D_{3}\rangle$. Contrary to this, the local magnetization exhibits at the higher magnetic field $B_{x}=5\,\mathrm{T}$ a more complex oscillatory dynamics with a clear amplitude modulation. Initially, the oscillations have a larger amplitude, but the amplitude decreases in time indicating a form of damped oscillation followed by a partial revival at later times. While the amplitude never reaches its initial value, this revival suggests a dynamic interference or rephasing effect potentially due to an external magnetic field. Although the dynamical oscillations of the sensor and readout spins are much more complex, the sensor and readout spins maintain a synchronized oscillations at small up to moderate magnetic fields across various initial states what highlights potential of the molecular compound Fe3 for indirect measurement of the magnetic field $B_{x}$ through the readout spin dynamics.
V.1 Sequential measurement

### v3

@@ III Quantum entanglement
The bipartite negativity [77] represents one of the most commonly used measures of the pairwise entanglement in the quantum information theory, which quantifies the amount of nonclassical correlations between two subsystems of a composite quantum system. Owing to the $\mathrm{C}_{3}$ symmetry of the trinuclear complex Fe3, all three reduced density matrices of the spin-5/2 Heisenberg triangle obtained after tracing out degrees of freedom of the third spin ${\rho}_{ij}=\mbox{Tr}_{k}\rho$ are completely identical ${\rho}_{12}={\rho}_{23}={\rho}_{13}\equiv{\rho}_{\mathrm{Bip}}$. Consequently, all three associated bipartite negativities ${N}_{12}={N}_{23}={N}_{13}\equiv{N}_{\mathrm{Bip}}$ will become identical as well. To quantify a strength of the bipartite entanglement between any pair of spin-5/2 magnetic ions iron(III) of the trinuclear complex Fe3 one can therefore utilize the formula put forward by Vidal and Werner [77], which is defined according to the Peres–Horodecki separability criterion [73, 74] as a sum of absolute values of negative eigenvalues of a partially transposed reduced density matrix $\rho_{\mathrm{Bip}}^{\mathrm{T}}$
@@
Here, $\lambda_{i}$ denote the eigenvalues of the reduced density matrix $\rho_{\mathrm{Bip}}^{\mathrm{T}}$ that is partially transposed with respect to one of two remaining spins. The reduced density matrix $\rho_{\mathrm{Bip}}$ and its transpose $\rho_{\mathrm{Bip}}^{\mathrm{T}}$ along with the eigenvalues $\lambda_{i}$ are obtained using exact numerical methods implemented in the QuTip package [78].
Figure 3: (a) Magnetic-field dependencies of the bipartite negativity ${N}_{\mathrm{Bip}}$ for a few different values of temperature; (b) temperature dependencies of the bipartite negativity ${N}_{\mathrm{Bip}}$ for a few selected values of the magnetic field. In both panels we present the theoretical prediction of the bipartite negativity ${N}_{\mathrm{Bip}}$ for any pair of high-spin iron(III) magnetic ions within the molecular complex Fe3 based on the spin-5/2 Heisenberg triangle, which is given by the Hamiltonian (1) with the coupling constant $J=12.56\,\mathrm{cm}^{-1}$ and the gyromagnetic factor $g=2.0$ as determined from the previous experiments [72].
@@
To determine whether all three spin-5/2 iron(III) magnetic ions in the trinuclear complex Fe3 are involved in multipartite quantum correlations, we further examine the tripartite entanglement within the respective spin-5/2 Heisenberg triangle. The tripartite quantum entanglement shared among all three spins is captured by the quantity genuine tripartite negativity ${N}_{\mathrm{Trip}}$, which is defined as the geometric mean of the bipartite negativities calculated for all three possible bipartitions of the spin-5/2 Heisenberg triangle [79]
@@
Figure 4: (a) Magnetic-field dependencies of the tripartite negativity ${N}_{\mathrm{Trip}}$ for a few different values of temperature; (b) temperature dependencies of the tripartite negativity ${N}_{\mathrm{Trip}}$ for a few selected values of the magnetic field. In both panels we present the theoretical prediction of the tripartite negativity ${N}_{\mathrm{Trip}}$ for any pair of high-spin iron(III) magnetic ions within the molecular complex Fe3 based on the spin-5/2 Heisenberg triangle, which is given by the Hamiltonian (1) with the coupling constant $J=12.56\,\mathrm{cm}^{-1}$ and the gyromagnetic factor $g=2.0$ as determined from the previous experiments [72].
@@
We now proceed with a theoretical examination of the genuine tripartite negativity ${N}_{\mathrm{Trip}}$ of the trinuclear complex Fe3 as shown in Fig. 4. We note that the QuTip package [78] is used here to numerically obtain the partially transposed density matrix $\rho_{\mathrm{ABC}}^{\mathrm{T}_{\mathrm{A}}}$ and its eigenvalues. It is evident from Fig. 4(a) that the behavior of the genuine tripartite negativity closely resembles at low temperatures that of its bipartite counterpart depicted in Fig. 3(a). Specifically, the genuine tripartite negativity ${N}_{\mathrm{Trip}}$ repeatedly exhibits unconventional step-like changes including a series of plateaus and sudden downturns as the magnetic field increases. As the temperature rises, this stepwise behavior of the genuine tripartite negativity ${N}_{\mathrm{Trip}}$ is gradually fading and transitioning into a smooth monotonic decrease with increasing magnetic field (see the lines for $T\gtrsim 10$ K in Fig. 4(a)). Additional insights can be drawn from the temperature dependencies of the genuine tripartite negativity in Fig. 4(b), which reveals that the molecular complex Fe3 displays a high degree of tripartite entanglement at low enough temperatures $T\lesssim 10$ K and magnetic fields $B\approx 10\,\mathrm{T}$. Under these conditions, the eigenstate $|1/2\rangle$ with the lowest possible spin value dominates the overall mixed state. Notably, the tripartite negativity gradually decreases as the temperature increases and eventually vanishes at a significantly higher threshold temperature $T\approx 70\,\mathrm{K}$ compared to the threshold temperature of the bipartite entanglement when exceeding it more than twice. This finding suggests that the trinuclear complex Fe3 remains subject to multipartite entanglement even at elevated temperatures and magnetic fields.
IV Quantum Sensing
@@
To adapt the quantum sensing protocol suggested in Ref. [54] for a real magnetic compound, let us consider the molecular complex Fe3 consisting of three interacting spin-5/2 iron(III) magnetic ions as a probe designed for sensing a local magnetic field $B_{x}$ acting on one iron(III) magnetic ion via measuring either of the remaining two iron(III) magnetic ions. The Hamiltonian of the spin-5/2 Heisenberg triangle with a local magnetic field acting only on a single spin, which may serve as the probe, can be updated to the following form:
@@
It has been demonstrated in Ref. [80] that the Dicke states $|D_{k}\rangle$ of three qudits can be straightforwardly obtained by applying $k$ times the total spin-raising operator $J_{+}$ on the basis state with the lowest total spin angular momentum $|-\frac{5}{2},-\frac{5}{2},-\frac{5}{2}\rangle$. More specifically, the Dicke states $|D_{k}\rangle$ of three spin-5/2 qudits can be defined as follows:
@@
To demonstrate how the molecular compound Fe3 can facilitate the implementation of a quantum sensing protocol described in Ref. [54], we specifically consider a local magnetic field acting on the first spin of the spin-5/2 Heisenberg triangle with the readout performed sequentially on either of the remaining two spins.
Figure 5: The time evolution of the local magnetization of the first and third spin of the spin-5/2 Heisenberg triangle with the coupling constant $J=12.56\,\mathrm{cm}^{-1}$ and the gyromagnetic factor $g=2.0$ adjusted for a theoretical modeling of the molecular complex Fe3 [72] by considering four different initial Dicke states: (a)–(b) $|D_{0}\rangle$; (c)–(d) $|D_{1}\rangle$; (e)–(f) $|D_{2}\rangle$; (g)–(h) $|D_{3}\rangle$. In the left panel we suppose a lower magnetic field $B_{x}=1\,\mathrm{T}$, while in the right panel we consider a higher magnetic field $B_{x}=5\,\mathrm{T}$. Both spins exhibit almost identical spin dynamics.
@@
After initializing, the state of the spin-5/2 Heisenberg triangle evolves under the influence of the Hamiltonian $\hat{H}^{\prime}$ and the dynamics of local magnetization $\langle S^{z}_{1}\rangle$ and $\langle S^{z}_{3}\rangle$ depends on the magnetic field. Fig. 5 illustrates the time evolution of the magnetization dynamics for different initial states under two different magnetic fields $B_{x}=1\,\mathrm{T}$ and $B_{x}=5\,\mathrm{T}$. Each panel compares the dynamical behavior of the readout spin $\langle S_{3}^{z}\rangle$ (pale lines) and the sensor spin $\langle S_{1}^{z}\rangle$ (blue dotted lines) for two different magnetic fields. Figs. 5(a) and 5(b) illustrate the time evolution of magnetization for the initial state $|D_{0}\rangle$ with character of the fully polarized ferromagnetic state. In Fig. 5(a), the sensor spin shows at a relatively small magnetic field $B_{x}=1\,\mathrm{T}$ oscillatory harmonic dependence and the readout spin closely follows this dependence. Fig. 5(b) shows the time evolution of the same quantities at the higher magnetic field $B_{x}=5\,\mathrm{T}$, for which the readout spin still mimics time evolution of the sensor spin but the oscillations are of higher frequency and less regular with gradually decreasing (and increasing) amplitude over the time. This scenario suggests more complex dynamical behavior as the magnetic field strengthens though the readout spin still continues to follow the sensor spin.
@@
Figs. 5(c)-5(h) depict the time evolution of magnetization for more complex initial Dicke states being subject to quantum superposition of states. At sufficiently small magnetic fields, the sensor and readout spins exhibit regular synchronized harmonic oscillations in time as demonstrated in Figs. 5(c), 5(e) and 5(g) for the particular case with $B_{x}=1\,\mathrm{T}$ and a few different Dicke states $|D_{1}\rangle$, $|D_{2}\rangle$, and $|D_{3}\rangle$. Contrary to this, the local magnetization exhibits at the higher magnetic field $B_{x}=5\,\mathrm{T}$ a more complex oscillatory dynamics with a clear amplitude modulation. Initially, the oscillations have a larger amplitude, but the amplitude decreases in time indicating a form of damped oscillation followed by a partial revival at later times. While the amplitude never reaches its initial value, this revival suggests a dynamic interference or rephasing effect potentially due to an external magnetic field. Although the dynamical oscillations of the sensor and readout spins are much more complex, the sensor and readout spins maintain a synchronized oscillations at small up to moderate magnetic fields across various initial states what highlights potential of the molecular compound Fe3 for indirect measurement of the magnetic field $B_{x}$ through the readout spin dynamics.
IV.1 Sequential measurement